A 122.74 g of FeSO4•xH2O contains 77.04 g of FeSO4, calculate the percent by mass of water in the hydrate and record the value of x in the short answer box.


1
Expert's answer
2021-07-18T01:27:39-0400

H2O= 122.74-77.04

45.7 g

% = 45.7/122.74×100 = 37.23%

Molar Mass of H2O = 18.015

x= 45.7/18.015

x= 2.54